Revision of Depreciation

Equipment with a cost of $612,200 has an estimated residual value of $55,700, has an estimated useful life of 35 years, and is depreciated by the straight-line method.

a. Determine the amount of the annual depreciation.

b. Determine the book value after 20 full years of use.


c. Assuming that at the start of the year 21 the remaining life is estimated to be 19 years and the residual value is estimated to be $58,600, determine the depreciation expense for each of the remaining 19 years.
Answer #1

a. Annual depreciation=($612200-$55700)/35=15900

b. Book value after 20 years=$612200-(15900*20)=294200

c. Deprexiation expense==294200-(58600-55700)/19=15331.57=15332